A total of 1700 in two mutual funds. One earned 6% the other earned 3% A profit of 60 dollars was earned. How much was invested in each mutual fund
Answer by mananth(16946)

Account I6.00% per annum ------------- Amount invested =x
Account II3.00% per annum ------------ Amount invested = y
1700
Interest----- 60

Account I 6.00% per annum ---x
Account II 3.00% per annum ---y

x + 1 y= 1700 -------------1

6.00% x + 3.00% y= 60
Multiply by 100
6 x + 3 y= 6000.00 --------2
Multiply (1) by -6
we get
-6 x -6 y= -10200.00
Add this to (2)
0 x -3 y= -4200
divide by -3
y = 1400
Account I 6.00% $ 300
Account II 3.00% $ 1400

CHECK
300 --------- 6.00% ------- 18.00
1400 ------------- 3.00% ------- 42.00
Total -------------------- 60.00

m.ananth@hotmail.ca